Through (9,-4), perpendicular to -5x-4y=-29
Answer by mananth(16949)

-5 x + -4 y = -29
Find the slope of this line

-4 y = 5 x -29
Divide by -4
y = -5/ 4 x + 29/4
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = -5/4

The slope of a line perpendicular to the above line will be the negative reciprocal 4/5
Because m1*m2 =-1
The slope of the required line will be 4/5

m= 4/5 ,point ( 9 , -4 )
Find b by plugging the values of m & the point in
y=mx+b
-4 = 36/ 5 + b
b= -56/5
m= 4/5
The required equation is y = 4/ 5 x -56/ 5
m.ananth@hotmail.ca